OpenStudy (anonymous):

Solve for the variable: (u/8)-12=6

OpenStudy (anonymous):

you gotta get the variable alone.

OpenStudy (anonymous):

Some of my answers are 122, 134, 144, 160

OpenStudy (anonymous):

..perfect so what u can do it try substituting the numbers in the varable to see if u get ur answer

OpenStudy (anonymous):

Wat

OpenStudy (anonymous):

sorry im just confused

OpenStudy (anonymous):

like replace u with one of the numbers and see if u get 6 ur answer is 144

OpenStudy (anonymous):

144/8 - 12 = 6 18 - 12 = 6 6 = 6

OpenStudy (anonymous):

Okay I have a couple more I need help on. Can u help me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve: 3 - 4 ( t+1) =7

OpenStudy (anonymous):

what r ur answers?

OpenStudy (anonymous):

-2, 0, 5/4, 3/4

OpenStudy (anonymous):

its one of the frctions

OpenStudy (anonymous):

But which one ? Is it 5/4 or 3/4

OpenStudy (anonymous):

idk.. im not good at doing math with fractions sorry :(

OpenStudy (anonymous):

Its okay and im not good at algerba

OpenStudy (anonymous):

im good at some.. but not all.

OpenStudy (anonymous):

Same here.

OpenStudy (anonymous):

I might need some help myself lol

OpenStudy (anonymous):

What do u need help on mayb I can help

OpenStudy (judygreeneyes):

3 - 4(t+1) = 7 First multiply the stuff inside the parentheses by 4: 3 - 4t -4 = 7 or, simplifying, -4t - 1 = 7 Now get the t on the left side, get the numbers all to the right side, by adding 1 to both sides: -4t = 8 Now to isolate t, divide both sides by -4: t = -2 Plug it back in to see if it is right: 3 - 4(-2+1) = 7 3 + 4 = 7 true!!